# In-Context Multi-Operator Learning with DeepOSets

### v1

@@ Abstract
In-context Learning (ICL) is the remarkable capability displayed by some machine learning models to learn from examples in a prompt, without any further weight updates. ICL had originally been thought to emerge from the self-attention mechanism in autoregressive transformer architectures. DeepOSets is a non-autoregressive, non-attention based neural architecture that combines set learning via the DeepSets architecture with operator learning via Deep Operator Networks (DeepONets). In a previous study, DeepOSets was shown to display ICL capabilities in supervised learning problems. In this paper, we show that the DeepOSets architecture, with the appropriate modifications, is a multi-operator in-context learner that can recover the solution operator of a new PDE, not seen during training, from example pairs of parameter and solution placed in a user prompt, without any weight updates. Furthermore, we show that DeepOSets is a universal uniform approximator over a class of continuous operators, which we believe is the first result of its kind in the literature of scientific machine learning. This means that a single DeepOSets architecture exists that approximates in-context any continuous operator in the class to any fixed desired degree accuracy, given an appropriate number of examples in the prompt. Experiments with Poisson and reaction-diffusion forward and inverse boundary-value problems demonstrate the ability of the proposed model to use in-context examples to predict accurately the solutions corresponding to parameter queries for PDEs not seen during training.
@@ 1 Introduction
In learning the solution operator of a parametric PDE from sample data, one is not interested in merely obtaining a single solution for a single PDE, but to learn the mapping between parameter and solution, so that new solutions corresponding to new values of the parameter can be obtained quickly and without further training. Furthermore, both the parameter and the solution are typically functions in infinite-dimensional spaces. For example, the problem might be to learn the space-varying fluid pressure (solution) corresponding to the space-varying permeability in a porous medium (parameter), where the physical process is governed by Darcy’s Equation. The training data in this case consists of pairs of parameter and corresponding solution, obtained from experiments or, more commonly, by solving Darcy’s equation using a high-fidelity solver. The trained operator allows quick exploration of the parameter space, which is unfeasible to do with expensive experiments or high-fidelity solves. The most well-known operator learners are DeepONets ludeeponet2021 and Fourier Neural Operators li2020fourier, which are both based on specific deep neural network architectures.
@@
A more recent development is foundation models for scientific problems, which is inspired by the success of of large language models. A foundation model can be trained on a large body of data and be adapted for new tasks, perhaps with no further training bommasani2021opportunities. In the context of PDEs, foundation models correspond to multiple operator (multi-operator) learning. The challenge could be to be able to train a model with data from diverse PDEs with diverse equation coefficients and boundary conditions, and then produce fast solutions for new PDEs on demand. In contrast to single operator learning, here a well-posedness question arises, since multiple solution operators acting on the same parameter function will produce different solutions. How is the model supposed to identify the correct operator? There are two basic methods to resolve this issue: the first, is to fine-tune the model on a small amount of data from the new PDE; the second is to use in-context learning (ICL), when the PDE is identified from example pairs in a prompt with no further training.
@@
In-Context Operator Networks (ICON) yang2023context and its subsequent variants yang2024pde; yang2023ICON-LM; zhang2025GenICON remain the only published approach to in-context learning of PDE solution operators (see Section 2 for other work related to multi-operator learning and foundation models for PDEs). ICON is a framework that adapts GPT-style in-context learning to PDE problems, by employing a decoder transformer to generate the solution from a sequential prompt followed by the query. ICON is thus an autoregressive, attention-based approach, with quadratic complexity in prompt length, due to the dot-product self-attention mechanism. This can introduce significant computational overhead and loss of accuracy with very long prompts. In addition, if the query changes, the prompt has to be processed sequentially again from the beginning (unless caching is used).
@@
In this paper, we propose an alternative approach to in-context learning of PDE solution operators, which processes the prompt non-autoregressively, in a fully parallel manner, without using self-attention, being therefore much faster during both training and inference time. This approach is based on DeepOSets, a neural architecture that combines set learning via the DeepSets architecture zaheer2017deep with operator learning via the previously mentioned DeepONets architecture. In DeepOSets, the DeepSets layer acts as an encoder of the prompt into a latent representation that is the input, along with the query, to the branch path of a DeepONet. The DeepSets layer confers linear complexity in the length of the prompt, in contrast to the quadratic complexity of self-attention, and a built-in permutation-invariance inductive bias to the order of the examples in the prompt.
@@
DeepOSets was shown in a previous publication to display ICL capabilities in regression problems chiu2024deeposets. Here, we introduce architectural modifications to DeepOSets that allow it to act as a multi-operator learner for parametric PDEs. This is accomplished by discretizing a variable number of solution and parameter function pairs, which are entered into the prompt and encoded by the DeepSets layer into latent representation. The latter is concatenated to the discretization of the query parameter function and entered directly into the branch path of the DeepONet.
@@
Crucially, we prove that this DeepOSets architecture approximates uniformly a compact set of continuous solution operators, which means that DeepOSets are not only universal approximators of single solution operators, like DeepONets, but they are universal approximators of multiple operators, in the sense that a single DeepOSets architecture exists that can approximate any operator in the class to the same fixed degree of accuracy. This guarantees that DeepOSets can learn in-context the solution operators of multiple PDEs.
@@
Experiments with Poisson and reaction-diffusion forward and inverse problems demonstrate that DeepOSets can accurately predict the solution to a new PDE with inverse or forward problem and unknown value of equation coefficients and boundary conditions learned in-context.
@@ 2 Related Work
MIONet is probably the first approach to attempt multi-input operator learning, by encoding several input functions via multiple branch network (and a shared trunk network) of a DeepONet, though it still learns a single operator jin2022mionet. POSEIDON is a large foundation model for PDEs based on a scalable Operator Transformer (scOT) with a vision transformer architecture, is pretrained on compressible Euler and incompressible Navier-Stokes equations and applied to 15 downstream tasks, showing better sample efficiency and accuracy than baselines herde2024poseidon. PROSE-PDE is a transformer-based multi-operator learning approach trained on one-dimensional time-dependent nonlinear PDEs that encodes both numerical and symbolic data types for forward and inverse PDE problems sun2025towards. LeMON is based on pretraining and fine-tuning for multi-operator learning, and introduces low-rank adaptation to reduce computational cost sun2024lemon. MT-DeepONet introduces modifications to the branch network of a DeepONet to accommodate various functional forms of parameterized coefficients and multiple geometries kumar2025synergistic. In-context learning for PDE solution operators was introduced by ICON yang2023context. The follow-up study on 1D scalar conservation laws provides a detailed methodology for applying ICON to temporally evolving PDEs yang2024pde, while ICON-LM integrates ICON concepts with language-model architectures yang2023ICON-LM. GenICON introduces a generative, probabilistic approach to ICON zhang2025GenICON.
3 The DeepOSets Architecture for In-Context Operator Learning
The DeepOSets model is a hybrid neural network comprising a DeepSets encoder and a DeepONet decoder, as seen in Figure 1. The DeepSets module encodes non-autoregressively the discretized parameter and solution function example pairs in the prompt. Each pair $(u_{i},G(u_{i}))$ is first mapped to an individual fixed-length embedding $\bf{h}_{i}$ by a shared multi-layer perceptron (MLP) denoted $\phi$. These intermediate representations are then combined by a pooling layer into the embedding vector $\bf{h}$. This introduces a permutation invariance inductive bias to the prompt examples, as long as the pooling layer is a permutation invariant operation, such as min, max, mean or set attention pooling. The DeepONet takes the concatenation of the embedding $\bf{h}$ and the discretized query parameter function $u_{q}$ at its branch path and the query coordinates $x_{q}$ at its trunk path, and outputs the predicted solution function $\hat{G}(u_{q})$ evaluated at $x_{q}$. The output is therefore entirely mesh-free.
4 Universality of DeepOSets for In-Context Operator Learning
In this section, we prove that DeepOSets are universal uniform approximators over a compact set of continuous operators, i.e., a single DeepOSets neural network exists that bounds the supremum of the error over the entire class to any desired fixed accuracy. This shows that a unique DeepOSets neural network could be learned that approximates in-context any continuous operator to the desired accuracy, given an appropriate number of examples in the prompt.
We first define the class of operators of interest. Let $X:=C(K_{1}),~Y:=C(K_{2})$, with $K_{1},K_{2}\subset\mathbb{R}^{d}$ compact subsets of some Euclidean space. Let $V\subset X$ be compact. Given this, consider the space of continuous operators (in the induced operator norm):
@@ 4 Universality of DeepOSets for In-Context Operator Learning
with norm given by
@@
We let $\mathcal{K}\subset\mathcal{O}$ to be compact subset of this space of operators. We will need to discretize functions in $X$ on different grids in $K_{1}$ and will also need to consider certain coverings of the set $V$. For this purpose, we introduce the following definition.
Definition 4.1 (Grid).
Let $\delta>0$ and $C>1$. A $(\delta,C)$-discretization of a compact metric space $K$ is a set of points $\textbf{x}=\{x_{1},\cdots,x_{n}\}$ such that
$|\{x_{i}\in B(x,\delta/2)\}|<C|\{x_{i}\in B(y,\delta)\}|$
for any $x,y\in K$.
Intuitively, a $(\delta,C)$-discretization is a $\delta$-net of $K$ which is not too unbalanced, in the sense that the density of points is relatively uniform. Let us first note that this definition is not vacuous.
@@
Let $K$ be a compact set in a metric space, $\delta>0$ and $C>1$. Then there always exists a $(\delta,C)$-discretization of $K$.
@@
Given a continuous function $u:K\rightarrow\mathbb{R}$ and a $(\delta,C)$-discretization $\textbf{x}=(x_{1},...,x_{n})$, we denote by
$u(\textbf{x}):=\left(u(x_{i})\right)_{i=1}^{n}\in\mathbb{R}^{n}$ (1)
for the function evaluated at the grid x. Given these definitions and notation, our main result is the following.

### v2

@@ Abstract
An important application of neural networks to scientific computing has been the learning of non-linear operators. In this framework, a neural network is trained to fit a non-linear map between two infinite dimensional spaces, for example, the solution operator of ordinary and partial differential equations. Recently, inspired by the discovery of in-context learning for large language models, an even more ambitious paradigm has been explored, called multi-operator learning. In this approach, a neural network is trained to learn many different operators at the same time. In order to evaluate one of the learned operators, the network is passed example inputs and outputs to disambiguate the desired operator. In this work, we provide a precise mathematical formulation of the multi-operator learning problem. In addition, we modify a simple efficient architecture, called DeepOSets, for multi-operator learning and prove its universality for multi-operator learning. Finally, we provide a comprehensive set of experiments that demonstrate the ability of DeepOSets to learn multiple operators corresponding to different initial-value and boundary-value differential equations and use in-context examples to predict accurately the solutions corresponding to queries and differential equations not seen during training. The main advantage of DeepOSets is its architectural simplicity, which allows the derivation of theoretical guarantees and training times that are in the order of minutes, in contrast to similar transformer-based alternatives that are empirically justified and require hours of training.
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Learning non-linear operators from data using neural networks has recently proved to be a foundational application of deep learning to scientific computing ((Lu et al., 2021; Li et al., 2020; Boullé and Townsend, 2024). In learning the solution operator of a parametric PDE from sample data, one is not interested in merely obtaining a single solution for a single PDE (as is typical in the popular Physics-informed neural network approach ((Raissi et al., 2019)), but instead learning the mapping between parameter and solution, so that new solutions corresponding to new parameter values can be obtained quickly and without further training. Furthermore, both the parameter and the solution are typically functions in infinite-dimensional spaces. For example, the problem might be to learn the space-varying fluid pressure (solution) corresponding to the space-varying permeability in a porous medium (parameter), where the physical process is governed by Darcy’s Equation. The training data in this case consists of pairs of parameters and corresponding solutions, obtained from experiments or, more commonly, by solving Darcy’s equation using a high-fidelity solver. The trained operator allows quick exploration of the parameter space, which is infeasible to do with expensive experiments or high-fidelity solvers. The most well-known operator learners are DeepONets ((Lu et al., 2021) and Fourier Neural Operators ((Li et al., 2020), which are both based on specific deep neural network architectures.
As in the case of standard multi-layer perceptrons, the theoretical foundations of this approach are a universal approximation theorem, originally due to Chen and Chen ((1995b), which show that a precursor to the DeepONet ((Lu et al., 2021) architecture is universal. Specifically, this means that DeepONet can in principle approximate to arbitrary accuracy any continuous non-linear map $F:K\rightarrow C(K_{2})$, where $K_{1}\subset\mathbb{R}^{d}$ and $K_{2}\subset\mathbb{R}^{n}$ are compact subsets and $K\subset C(K_{1})$ is a compact subset of $C(K_{1})$.
@@
A more recent development is foundation models for scientific problems, which is inspired by the success of large language models. A foundation model can be trained on a large body of data and be adapted for new tasks, perhaps with no further training ((Bommasani, 2021). In the context of differential equations, foundation models correspond to multiple operator (multi-operator) learning. The challenge here could be to be able to train a model with data from diverse differential equations with diverse equation coefficients and boundary conditions, and then produce fast solutions for new PDEs on demand. In contrast to single operator learning, here a well-posedness question arises, since multiple solution operators acting on the same parameter function will produce different solutions. How is the model supposed to identify the correct operator? There are two basic methods to resolve this issue: the first, is to fine-tune the model on a small amount of data from the new equation; the second is to use in-context learning (ICL), when the equation is identified from example pairs in a prompt with no further training.
@@
In-Context Operator Networks (ICON) ((Yang et al., 2023a) and its subsequent variants ((Yang and Osher, 2024; Yang et al., 2023c; Zhang et al., 2025) is the best-known approach to in-context learning of PDE solution operators (see Section 2 for other work related to multi-operator learning and foundation models for PDEs). ICON is a framework that adapts GPT-style in-context learning to differential equation problems, by employing a decoder transformer to generate the solution from a sequential prompt followed by the query. ICON is thus an attention-based approach, with quadratic complexity in prompt length, due to the dot-product self-attention mechanism. This can introduce significant computational overhead and loss of accuracy with very long prompts. In addition, if the query changes, the prompt has to be processed again (unless caching is used).
In addition, a theoretical study of in-context multi-operator learning remains incomplete. In particular, it is not known whether existing architectures such as ICON are universal approximators. One key difficulty is that the lack of a precise mathematical formulation of the multi-operator learning problem, i.e., what does it mean to simultaneously approximate multiple operators at the same time? In this paper, we provide a precise mathematical formulation of the in-context multi-operator learning problem, which is formulated as learning a compact set of operators between two Banach spaces, instead of just a single operator.
@@
Further, we propose an alternative approach to in-context learning of PDE solution operators, which processes the prompt non-autoregressively, in a fully parallel manner, without using self-attention, being therefore much faster during both training and inference time. This approach is based on DeepOSets, a neural architecture that combines set learning via the DeepSets architecture ((Zaheer et al., 2017) with operator learning via the previously mentioned DeepONets architecture. In DeepOSets, the DeepSets layer acts as an encoder of the prompt into a latent representation that is the input, along with the query, to the branch path of a DeepONet. The DeepSets layer confers linear complexity in the length of the prompt, in contrast to the quadratic complexity of self-attention, and a built-in permutation-invariance inductive bias to the order of the examples in the prompt.
@@
DeepOSets has been shown to display ICL capabilities in regression problems ((Chiu et al., 2024). Here, we introduce architectural modifications to DeepOSets that allow it to act as a multi-operator learner for parametric differential equations. This is accomplished by discretizing a variable number of solution and parameter function pairs, which are entered into the prompt and encoded by the DeepSets layer into latent representation. The latter is concatenated to the discretization of the query parameter function and entered directly into the branch path of the DeepONet. A detailed description of our architecture can be found in Section 4 and Figure 1.
@@
Crucially, we prove that this DeepOSets architecture is universal for the in-context multi-operator learning problem. Specifically, we prove that it can approximate uniformly any compact set of continuous solution operators. This means that DeepOSets are not only universal approximators of single solution operators, like DeepONet, but they are universal approximators of multiple operators, in the sense that a single DeepOSets architecture exists that can approximate any operator in the class to the same fixed degree of accuracy. This guarantees that DeepOSets can learn in-context the solution operators of multiple differential equations.
@@
Finally, experiments with Poisson and reaction-diffusion forward and inverse problems demonstrate that DeepOSets can accurately predict the solution to a new differential equation with inverse or forward versions and unknown value of equation coefficients and boundary conditions learned in-context. Training times are in the order of minutes, as compared to hours for similar transformer-based alternatives.
1.1 Our Contributions
To summarize, our main contributions are:
  1. 1.
    We provide a precise mathematical formulation of the in-context multi-operator learning problem and universal approximation problem in Section 3 and Definition 3.3, respectively.
  2. 2.
    We prove that a variant of DeepOSets, defined in Section 4, is universal for in-context multi-operator learning within Section 5 in Theorem 5.1. This is to the best of our knowledge the first universal approximation result for in-context multi-operator learning.
  3. 3.
    We demonstrate in Section 6 that DeepOSets is capable of jointly learning multiple operators and exhibits robust generalization to support sets larger than those encountered during training. Furthermore, training is very efficient, taking only minutes on a consumer GPU.
@@ 2 Related Work
The universal approximation theorem for multi-layer perceptrons, which forms the basis for more sophisticated universal approximation theorems, was proved in Cybenko ((1989) and Hornik ((1991). Universal approximation of operators using neural networks was first proved in Chen and Chen ((1995b), following previous results from Chen and Chen ((1993, 1995a). This result was used to design the DeepONet operator learning architecture ((Lu et al., 2021). Subsequently, various universal architectures for operator learning have been proposed, a notable example being Fourier Neural Operators ((Li et al., 2020). These architectures typically evaluate the input function on a fixed grid, and an architecture compatible with variable input grids was proposed and analyzed in Prasthofer et al. ((2022).
In-context learning (ICL) is a phenomenon first observed in large language models ((Brown et al., 2020; Tsimpoukelli et al., 2021; Liu et al., 2023; Garg et al., 2022), whereby these models are able to learn new tasks without fine-tuning from examples presented to them in the prompt. This has sparked substantial theoretical research effort into the mechanism underlying this phenomenon. Transformers have been theoretically shown to exhibit ICL on a variety of linear models, including least squares, ridge regression, Lasso, generalized linear models and linear inverse problems ((Bai et al., 2023; Zhang et al., 2024; Von Oswald et al., 2023). Generalization error for transformer-based ICL for data lying on a low-dimensional manifold has been studied in ((Shen et al., 2025).
@@
In terms of applying ICL to operator learning, MIONet is to the best of our knowledge the first approach to attempt multi-input operator learning, by encoding several input functions via multiple branch network (and a shared trunk network) of a DeepONet, though it still learns a single operator ((Jin et al., 2022). POSEIDON is a large foundation model for PDEs based on a scalable Operator Transformer (scOT) with a vision transformer architecture, is pretrained on compressible Euler and incompressible Navier-Stokes equations and applied to 15 downstream tasks, showing better sample efficiency and accuracy than baselines ((Herde et al., 2024). PROSE is a transformer-based multi-operator learning approach trained on one-dimensional time-dependent nonlinear PDEs that encodes both numerical and symbolic data types for forward and inverse PDE problems ((Sun et al., 2025). PROSE is used to build a foundation model for fluid dynamics problems in Liu et al. ((2024). Recently, the PROSE approach was extended to include physics-informed losses ((Zhu et al., 2025). LeMON is based on pretraining and fine-tuning for multi-operator learning, and introduces low-rank adaptation to reduce computational cost ((Sun et al., 2024). MT-DeepONet introduces modifications to the branch network of a DeepONet to accommodate various functional forms of parameterized coefficients and multiple geometries ((Kumar et al., 2025). In-context learning for differential equation solution operators was introduced by ICON ((Yang et al., 2023a), with a follow-up study on 1D scalar conservation laws providing a detailed methodology for applying ICON to time-evolution PDEs ((Yang and Osher, 2024). ICON-LM integrates ICON concepts with language-model architectures ((Yang et al., 2023c). GenICON introduces a generative, probabilistic approach to ICON ((Zhang et al., 2025). In-context learning of linear PDEs using transformers has also been studied in ((Cole et al., 2024).
3 Mathematical Formulation of In-context Operator Learning
In this section, we give a precise mathematical formulation of the in-context multi-operator learning problem. Based upon this, we will be able to argue whether a given architecture is universal for this problem.
Let $K_{1},K_{2}\subset\mathbb{R}^{d}$ be compact subsets of some Euclidean space. We will consider continuous maps between the Banach spaces $X:=C(K_{1})$ and $Y:=C(K_{2})$, whose domain is a compact subset $V\subset X$. Note that the dimensionality of $K_{1}$ and $K_{2}$ need not be the same in general, but to simplify the presentation we restrict to this case in the following. The minor adjustments required to handle the more general case where $K_{1}\subset\mathbb{R}^{d_{1}}$ and $K_{2}\subset\mathbb{R}^{d_{2}}$ are left to the reader.
In traditional operator learning, a continuous function $G:V\rightarrow Y$ is approximated using a neural network. The input function $u\in V$ is discretized at a fixed set of points $x_{1},...,x_{k}\in K_{1}$, and the values $u(x_{1}),...,u(x_{k})$ together with an evaluation point $y\in K_{2}$ are passed to a neural network whose output approximates $G(u)(y)$. Universality in this context means that for any $G$ and $\epsilon>0$, there exists a grid $\textbf{x}:=\{x_{1},...,x_{k}\}\in K_{1}$ (depending on $G$ and $\epsilon$) and a neural network $G_{\epsilon}$ such that
$\left|G_{\epsilon}(u(\textbf{x}),y)-G(u)(y)\right|<\varepsilon$
for any $u\in V$ and $y\in K_{2}$ ((Chen and Chen, 1995b; Lu et al., 2021). Here we denote by $u(\textbf{x}):=\left(u(x_{i})\right)_{i=1}^{k}\in\mathbb{R}^{k}$ the function evaluated at the grid x.
Instead of learning a single operator $G:V\rightarrow Y$, in-context multi-operator learning considers an entire class of operators. To formulate this, we consider the space of continuous operators
@@ 3 Mathematical Formulation of In-context Operator Learning
equipped with the induced operator norm given by
@@
We let $\mathcal{K}\subset\mathcal{O}$ be a compact subset of this space of operators. $\mathcal{K}$ is the class of operators that we wish to learn. In order to disambiguate the desired operator $G\in\mathcal{K}$, we pass the network a sequence of in-context examples $(u_{1},G(u_{1})),...,(u_{m},G(u_{m}))$ in addition to the input function $u\in K$ and the evaluation point $y\in K_{2}$. All of the functions $_{i}, uu$, and $G(u_{i})$ are discretized on fixed grids $\textbf{x}=(x_{1},...,x_{k})$ and $\textbf{y}=(y_{1},...,y_{n})$ in $K_{1}$ and $K_{2}$, respectively. We remark that one could additionally consider variable grids in $K_{1}$ and $K_{2}$, but we leave this to future work.
Although the grids x and y are fixed, we want to allow the in-context examples to be given for arbitrary functions $u_{1},...,u_{m}\in V$ (even the number of functions $m$ should also be arbitrary). Of course, we need to make some assumption which guarantees that the $u_{i}$ sufficiently cover the domain $V$ so that the desired operator $G\in\mathcal{K}$ can be determined to the desired accuracy from the outputs $G(u_{i})$. For technical reasons, we also need to choose the $u_{i}$ to be sufficiently uniformly distributed within $V$. This is captured in the following definition.
Definition 3.1 ($(\delta,C)$-discretization).
Let $\delta>0$ and $C>1$. A $(\delta,C)$-discretization of a compact metric space $V$ is a set of points $\textbf{u}=\{u_{1},\cdots,u_{m}\}$ such that
$|\{u_{i}\in B(u,\delta/2)\}|<C|\{u_{i}\in B(v,\delta)\}|\,$,
for any $u,v\in V$.
Intuitively, a $(\delta,C)$-discretization is a $\delta$-net of $V$ which is not too unbalanced, in the sense that the density of points is relatively uniform. Let us first note that this definition is not vacuous.
@@
Let $V$ be a compact set in a metric space, $\delta>0$ and $C>1$. Then there always exists a $(\delta,C)$-discretization of $V$.
@@
We can now formulate the concept of universality for in-context multi-operator learning.
Definition 3.3 (Mathematical formulation of In-Context Operator Learning).
We define universality for in-context operator learning to mean that for any $\mathcal{K}\subset\mathcal{O}$, any $\epsilon>0$ and $C>1$, there exist grids $\textbf{x}=(x_{1},...,x_{k})$ and $\textbf{y}=(y_{1},...,y_{n})$ in $K_{1}$ and $K_{2}$, respectively, a $\delta>0$ and a network $K_{\epsilon}$ such that
$|K_{\epsilon}(u_{1}(\textbf{x}),G(u_{1})(\textbf{y}),...,u_{m}(\textbf{x}),G(u_{m})(\textbf{y}),u(\textbf{x}),y)-G(u)(y)|<\epsilon\,$, (1)
for any $G\in\mathcal{K}$, $u\in V$, $y\in K_{2}$, and $(\delta,C)$-discretization $\textbf{u}=(u_{1},...,u_{m})$ of $V$. In order words, any operator of $\mathcal{K}$ can be approximated to arbitrary accuracy provided that the in-context examples are sufficiently dense in $V$.
Note that the network architecture should be able to process an arbitrary, i.e., not a priori specified, number of in-context examples. This can be achieved using a transformer architecture ((Vaswani et al., 2017), for example. In the next section, we describe a simple architecture based upon Deep Sets ((Zaheer et al., 2017), called DeepOSets ((Chiu et al., 2024) for this purpose.
4 The DeepOSets Architecture for In-Context Operator Learning
The DeepOSets model is a hybrid neural network comprising of a DeepSets encoder and a DeepONet decoder, as seen in Figure 1.
The DeepSets module encodes the discretized parameter and solution function example pairs in the prompt. Each pair $(u_{i},G(u_{i}))$ is first mapped to an individual fixed-length embedding $\bf{h}_{i}$ by a shared multi-layer perceptron (MLP) denoted $\Phi$. These intermediate representations are then combined by a pooling layer into the embedding vector $\bf{h}$. This introduces a permutation invariance inductive bias to the prompt examples, as long as the pooling layer is a permutation invariant operation, such as min, max, mean or set attention pooling.
The DeepONet takes the concatenation of the embedding $\bf{h}$ and the discretized query parameter function $u_{q}$ at its branch path and the query coordinates $x_{q}$ at its trunk path, and outputs the predicted solution function $\hat{G}(u_{q})$ evaluated at $x_{q}$. The output is therefore entirely mesh-free. We denote the branch network as $\rho$, and the trunk network as $\Psi$. With this notation, the output of the DeepOSets architecture using mean pooling, given input $x_{q}$, can be mathematically expressed as
$\hat{G}(u_{q})(x_{q}):=\Psi(x_{q})\cdot\rho\left(u_{q}(\textbf{x}),\frac{1}{m}\sum_{i=1}^{m}\Phi(u_{i}(\textbf{x}),G(u_{i})(\textbf{y}))\right).$ (2)
5 Universality of DeepOSets for In-Context Operator Learning
In this section, we prove that DeepOSets is a universal uniform approximator for any compact set of continuous operators in the sense of Definition 3.3, i.e., a single DeepOSets neural network exists that bounds the supremum of the error over the entire class to any desired fixed accuracy. This shows that a unique DeepOSets neural network could be learned that approximates in-context any continuous operator to the desired accuracy, given an appropriate number of examples in the prompt.